# VDGD: Mitigating LVLM Hallucinations in Cognitive Prompts by Bridging the Visual Perception Gap

### v1

@@ 5.2 Disentangling Visual Recognition from Other Cognitive Skills
A fundamental goal of alignment fine-tuning is to enable an LVLM to master implicit visual perception. As discussed in Section 3, when faced with an information-seeking or reasoning prompt, an LVLM is expected to respond directly to the prompt and to have implicitly comprehended the accompanying image with respect to the prompt instruction. To determine whether LVLM hallucinations stem from the limitations of the base LLM or from alignment issues, we aim to isolate visual recognition from other cognitive skills and evaluate them independently. To achieve this, we ask GPT-4 to rewrite the instructions in prompts from various visual information-seeking and reasoning benchmarks, such that they can be answered without the accompanying image. We show examples in Appendix H.6. Subsequently, we prompt LVLMs with these modified prompts without the image. Next, we prompt the LVLM to separately generate a description of the input image and evaluate this description for its faithfulness and detailedness (i.e., how well the image description can aid in a factual response to the original prompt). Evaluation prompts are provided in Appendix D. Fig. 7 (left) compares the performance of text-only prompts (denoted with -t in the key) with their original image-text counterpart across various LVLMs and Fig. 7 (right) shows their VR scores. As we see, in all cases, LVLMs fare well in VI, and are able to provide better factual responses than with the original prompt together with the input image. Thus, we conclude that while LVLMs demonstrate proficiency in visual recognition and have sufficient knowledge and reasoning abilities to generate factual responses, they encounter difficulties in employing these skills when responding to prompts that require their simultaneous application. This underscores a significant gap in alignment via lack of visual perception: LVLMs can recognize visual elements in images but fail to fully perceive them to accurately link what they recognize to their internal knowledge for reasoning or information extraction. We hypothesize a possible source of this issue to be the fact that most fine-tuning (or alignment) datasets for LVLMs consist of image-description pairs, which makes them struggle with accurate visual perception for non-description prompts.
@@ 5.3 Uncertainty in the Logit Space
Models Dataset k Variance Range Avg.
LLaVA-v1 AMBER 4.19 0.05 0.31 0.23
LLaVA-v1.5 AMBER 3.91 0.07 0.39 0.29
CogVLM AMBER 3.39 0.09 0.46 0.21
@@
LLaVA-v1.5 MMMU 3.53 0.09 0.41 0.23
CogVLM MMMU 3.16 0.19 0.47 0.25
To investigate the origins of hallucinations, we analyze the logit space of hallucinated tokens. Specifically, after applying the softmax operation to the logits, we rank them by their probability scores. We then use the elbow method (detailed in Appendix H.5) to truncate the distribution, obtaining the top-k tokens along with their corresponding probabilities. This truncation approach, based on the elbow method, effectively identifies the tokens most likely to be selected during multinomial sampling. Table 1 shows several statistics of the top-k probabilities for hallucinated tokens (we only consider the first token of each word). As evident from the value of k, standard deviation, and the average, the top-k probability space for hallucinated tokens is dominated by a set of low and equally confident tokens. Thus, each of these tokens has almost a similar chance of being sampled with multi-nomial sampling.
@@
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Our analysis of disentangling VR from other cognitive skills reveals that LVLMs generally demonstrate accurate VR and possess the necessary skills to respond accurately. However, they struggle to effectively link they recognize to their internal knowledge when responding directly, which inhibits the accurate execution of these skills. This issue underlines what we identify as the visual perception gap.
@@ 6 Visual Description Grounding Decoding (VDGD)
Figure 8: Illustration of VDGD Hallucination Mitigation.
To bridge the visual perception gap and reduce hallucinations, we propose Visual Description Grounding Decoding (VDGD), a simple, effective, and training-free hallucination mitigation technique. Specifically, during the decoding process, VDGD first adds the image description as a prefix to the text instruction of the prompt and then grounds token generation to the description of the input image by selecting a token that deviates the least from the description. We measure token deviation from the description with a novel formulation that employs Kullback–Leibler divergence at its core. The description is generated by the LVLM itself. VDGD operates on 2 core hypotheses: (1) Since LVLMs rely on language priors for cognitive prompts (Section 5.1), adding the image description as a prefix to the instruction can lead to improved visual perception by explicitly conditioning the response on the description. (2) Re-scoring the equally and low-confidence tokens (Section 1) based on deviation from the description can increase the confidence of the accurate token from several low-confidence tokens xu2023look .
@@ 6.1 Methodology
Fig. 8 illustrates the VDGD methodology. For an input prompt, we first employ the LVLM to generate a description of the image accompanying the prompt (see Appendix D for the prompt). Next, we concatenate the generated description as a prefix to the original prompt.
@@
Here, $p_{VLM}(x_{i}\mid x_{<i})$ denotes the probability distribution for the next token, $x_{i}$, given the previous tokens. Inspired by our findings in Section 5.3, we first apply a plausibility constraint to truncate the vocabulary space of the next token $x_{i}$ li-etal-2023-contrastive . Essentially, this step cuts off tokens from the vocabulary $\mathcal{V}$ below the elbow and keeps only the top-k plausible tokens. This is done as follows:
$p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\begin{cases}p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right),&\text{ if }p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\geq\alpha\max_{w}p_{\operatorname{VLM}}\left(w\mid x_{<i}\right)\leq\alpha\\ -\infty,&\text{ Otherwise }\end{cases}$ (2)
@@
6.2 VaLLu Benchmark
@@
Automatic evaluation of LVLMs, with powerful closed-source models, on multiple large benchmarks can be prohibitively expensive. Thus, we propose VaLLu, a benchmark that we build by amalgamating existing benchmarks. VaLLu comprises a total of 1500 instances, with 554 sourced from Oven, 535 from MMMU, 218 from MMC, 74 from MathVista, 60 from HallusionBench, 53 from MATH-Vision, and 14 from MME fu2023mme . We only include instances with instructions that require open-ended generations and do not include any binary Yes/No or Multi-choice questions. We control the size to ensure the evaluation is affordable while keeping the diversity of tasks and topics for comprehensive analysis. VaLLu is manually filtered (we found various noisy examples in existing benchmarks that do not have a definite ground-truth response and thus make evaluation difficult, we discuss this further in Appendix H.3), annotated (we annotate each instance for the type of skill required by the LVLM to generate a response, we illustrate this in Fig. 9) and answered (we accompany each instance with its ground-truth response).
@@ 6.3 Results and Analysis
Experimental Setup and Baselines. We compare VDGD against VCD, OPERA, Woodpecker, LRV, and LURE mitigation techniques. Additionally, we compare two vanilla decoding methods (without any additional mitigation techniques): Vanilla-greedy with vanilla greedy decoding and Vanilla-sampling with multinomial sampling-based decoding (top-p=0.5 and temperature=0.7). We employ greedy decoding for all methods as we find no difference in performance on sampling. We generate image descriptions for VDGD with the VCD mitigation technique and ablate the same.
Quantitative results. Table 2 compares the performance of VDGD with other baselines on the VaLLu Benchmark. We report factuality scores averaged across 3 runs for 3 random seeds. As we clearly see, VDGD not only outperforms all baselines but is also the only method that alleviates and improves performance on VaLLu. Results on other full benchmarks, including MMU, MathVista, and MATH-Vision, are provided in Appendix I.5, together with scores for other aspects beyond factuality where we show that VDGD does not compromise the depth and expressivity of responses. Removing KL divergence or just adding the description prefix results in a drop in performance.
Qualitative examples.Fig 10 compares responses from vanilla greedy decoding with VDGD. Grounding response generation results in more factual reasoning and information extraction.
Figure 10: Illustration of responses generated with VDGD compared with greedy decoding. More in Sec. H.2.
Methodology LLaVA-v1 LLaVA-v1.5 LLaVA-v1.6 mPLUG-Owl2 InternLM-X CogVLM
Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82
Vanilla-sampling 1.86 2.01 2.64 2.18 2.70 2.83
VCD 1.47 1.55 1.80 1.80 2.32 2.63
OPERA 2.04 2.05 2.62 2.26 2.65 2.85
Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91
LRV 2.03 2.10 2.64 2.24 2.64 2.78
LURE 1.98 2.03 2.65 2.19 2.59 2.88
@@
VDGD (-) VCD $2.08_{\pm 0.09}$ $2.43_{\pm 0.15}$ $3.01_{\pm 0.08}$ $2.54_{\pm 0.07}$ $3.26_{\pm 0.12}$ $2.95_{\pm 0.06}$
VDGD (-) KLD $2.05_{\pm 0.10}$ $2.30_{\pm 0.10}$ $2.78_{\pm 0.12}$ $2.37_{\pm 0.06}$ $3.01_{\pm 0.02}$ $2.87_{\pm 0.08}$
Table 2: Comparison of VDGD with other hallucination mitigation techniques. All results are on the proposed VaLLu benchmark for the "Factuality" metric. Results on other benchmarks can be seen in Appendix I.5.

### v2

@@ 4.2 Disentangling Visual Recognition from other Cognitive Skills
A key objective of alignment fine-tuning is to equip LVLMs with implicit visual perception skills. As outlined in Section 3, LVLMs should interpret images in context and respond accurately to information-seeking or reasoning prompts. To determine if hallucinations arise from the base LLM’s limitations or alignment gaps, we aim to isolate visual recognition from other cognitive skills and evaluate them separately. We use GPT-4 to rephrase prompts from visual information-seeking and reasoning benchmarks, modifying them so they can be answered without the image (examples in Appendix K.7 and prompt used in Appendix G).
[…]
[trim= 0cm 0.5cm 0cm 0.5cm, width=]lm_radar_combined.pdf
@@
LVLMs are first evaluated on these text-only prompts and then on generating independent image descriptions (evaluation prompt in Appendix G). Figure 7 (left) compares the performance of LVLMs on the text-only prompts (marked with -t) against their original image-text versions, while Figure 7 (right) shows VR scores. Our results indicate that LVLMs excel in visual recognition and perform better on factual responses when using text-only prompts than with combined image-text inputs. This suggests that LVLMs have the necessary knowledge and reasoning skills but struggle to integrate these with visual perception in prompts requiring both. We identify this as a perception gap: LVLMs can recognize visual elements but fail to fully perceive them in a way that connects recognition to internal knowledge, which is essential for reasoning or information extraction. This challenge likely arises due to several factors: (i) training datasets that are largely composed of image-description pairs, (ii) alignment datasets that contain strong visual cues within the natural-language prompts.
@@ 4.3 Uncertainty in the Logit Space
[…]
Table 1: Post-truncation probability statistics using the elbow method. All values are in the format: non-/hallucinated averaged across all tokens in the dataset.
To investigate the origins of hallucinations, we analyze the logit space of hallucinated tokens. Specifically, after applying the softmax operation to the logits, we rank them by their probability scores. We then use the elbow method (detailed in Appendix K.6) to truncate the distribution, obtaining the top-k tokens along with their corresponding probabilities. This truncation approach, based on the elbow method, effectively identifies the tokens most likely to be selected during multinomial sampling. Table 1 shows several statistics of the top-k probabilities for hallucinated tokens (we only consider the first token of each word). As evident from the value of k, standard deviation, and the average, the top-k probability space for hallucinated tokens is dominated by a set of low and equally confident tokens. Thus, each of these tokens has almost a similar chance of being sampled with multi-nomial sampling.
@@
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: LVLMs generally demonstrate accurate recognition and cognition skills to respond and reason accurately. However, they struggle to effectively link they recognize to their internal knowledge during the reasoning process, which results in inaccurate responses. This issue underlines what we identify as the visual perception gap.
@@ 5 Visual Description Grounding Decoding (VDGD)
[…]
[trim= 0cm 0.25cm 0cm 0.9cm, width=0.95]vllm-kl4.pdf
Figure 8: Illustration of our proposed VDGD method.
Main Motivation. To bridge the visual perception gap and reduce hallucinations for cognitive prompts, we propose Visual Description Grounding Decoding (VDGD), a simple, effective, and training-free hallucination mitigation technique. During decoding, VDGD first appends the image description (generated by the LVLM itself) as a prefix to the prompt’s text instruction, then grounds token generation by selecting the token that deviates the least from the description. Token deviation is measured using a novel formulation based on Kullback–Leibler divergence of the token with the description generated by the LVLM itself. Much like humans, who often write down their key observations in an image and refer to them while solving complex reasoning tasks, we hypothesize that grounding the LVLM’s responses in a detailed visual description can significantly enhance its reasoning process. Additionally, re-scoring the equally and low-confidence tokens (discussed in Section 1) based on deviation from the description can increase the confidence of the accurate token from several low-confidence tokens (Xu et al., 2023).
@@ 5.1 Methodology
Fig. 8 illustrates the VDGD methodology. For an input prompt, we first employ the LVLM to generate a description of the image accompanying the prompt (see Appendix G for the prompt). Next, we concatenate the generated description as a prefix to the original prompt.
@@
Here, $p_{VLM}(x_{i}\mid x_{<i})$ denotes the probability distribution for the next token, $x_{i}$, given the previous tokens. Inspired by our findings in Section 4.3, we first apply a plausibility constraint to truncate the vocabulary space of the next token $x_{i}$ (Li et al., 2023). Essentially, this step cuts off tokens from the vocabulary $\mathcal{V}$ below the elbow and keeps only the top-k plausible tokens. This is done as follows:
$p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\begin{cases}p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right),&\text{ if }p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\geq\alpha\max_{w}p_{\operatorname{VLM}}\left(w\mid x_{<i}\right)\\ -\infty,&\text{ Otherwise }\end{cases}$ (2)
@@
[…]
! Benchmark Baseline LLaVA-v1 LLaVA-1.5 LLaVA-1.6 mPLUG-Owl2 InternLM-X CogVLM CogVLM2 Qwen2-VL MMMU Vanilla-greedy 1.26 1.35 1.42 1.40 2.01 1.66 2.08 2.39 Vanilla-sampling 1.27 1.44 1.40 1.41 2.05 1.64 2.10 2.35 VDD 1.34 1.52 1.44 1.53 2.22 1.68 2.14 2.64 OPERA 1.30 1.43 1.57 1.64 2.25 1.62 2.21 2.44 Woodpecker 1.32 1.44 1.63 1.61 2.12 1.65 2.14 2.56 LRV 1.29 1.49 1.61 1.58 2.08 1.59 2.18 2.48 LURE 1.31 1.47 1.60 1.64 2.27 1.66 2.27 2.71 PAI 1.42 1.39 1.44 1.56 2.23 1.65 2.29 2.64 HALC 1.40 1.54 1.63 1.65 2.15 1.67 2.24 2.69 VDGD (ours) 1.49 (+18%) 1.62 (+20%) 1.75 (+23%) 1.72 (+23%) 2.39 (+19%) 1.71 (+3%) 2.47 (+19%) 2.91 (+22%) MathVista Vanilla-greedy 1.56 1.65 2.00 1.92 2.56 2.15 2.45 2.97 Vanilla-sampling 1.54 1.68 1.91 1.94 2.52 2.19 2.46 2.95 VDD 1.67 1.68 2.07 2.11 2.59 2.53 2.81 3.19 OPERA 1.64 1.83 2.04 2.04 2.62 2.30 2.53 3.22 Woodpecker 1.72 2.10 2.19 2.13 2.43 2.43 2.59 3.13 LRV 1.68 1.68 2.17 1.99 2.61 2.20 2.62 3.30 LURE 1.59 2.03 2.21 2.07 2.37 2.45 2.49 3.13 PAI 1.68 1.85 1.97 2.08 2.63 2.20 2.60 3.39 HALC 1.71 2.05 2.17 1.97 2.54 2.24 2.77 3.38 VDGD (ours) 1.84 (+18%) 2.19 (+33%) 2.44 (+22%) 2.24 (+17%) 2.88 (+13%) 2.62 (+22%) 2.93 (+20%) 3.59 (+21%) MMBench Vanilla-greedy 2.67 2.89 3.11 2.98 3.58 3.19 3.71 3.93 Vanilla-sampling 2.69 2.92 3.12 3.02 3.59 3.22 3.68 3.95 VDD 2.80 3.03 3.09 3.22 3.93 3.28 3.81 4.20 OPERA 2.81 2.94 3.17 3.10 3.95 3.22 3.93 4.05 Woodpecker 2.83 2.96 3.14 3.07 3.84 3.33 3.83 4.19 LRV 2.77 2.99 3.13 3.23 3.84 3.25 3.77 4.05 LURE 2.72 2.97 3.13 3.26 3.72 3.21 3.82 4.16 PAI 2.75 3.00 3.04 3.16 3.94 3.35 3.97 4.05 HALC 2.73 2.93 3.02 3.18 3.81 3.27 4.04 4.08 VDGD (ours) 2.93 (+10%) 3.12 (+8%) 3.26 (+5%) 3.32 (+11%) 4.08 (+14%) 3.42 (+7%) 4.15 (+12%) 4.31 (+10%) MME Vanilla-greedy 3.32 3.54 3.65 3.49 3.75 3.57 3.64 3.86 Vanilla-sampling 3.34 3.53 3.62 3.48 3.77 3.58 3.66 3.85 VDD 3.46 3.61 3.77 3.59 3.96 3.67 3.95 4.21 OPERA 3.42 3.59 3.82 3.54 3.93 3.60 3.88 4.09 Woodpecker 3.37 3.55 3.76 3.63 3.82 3.59 3.92 4.15 LRV 3.43 3.63 3.78 3.52 3.95 3.61 3.96 3.99 LURE 3.42 3.62 3.87 3.67 3.93 3.72 3.77 4.10 PAI 3.38 3.56 3.81 3.55 3.89 3.69 3.82 4.04 HALC 3.47 3.58 3.83 3.71 4.01 3.59 3.80 4.14 VDGD (ours) 3.59 (+8%) 3.70 (+5%) 3.99 (+9%) 3.82 (+9%) 4.12 (+10%) 3.79 (+6%) 4.09 (+12%) 4.34 (+12%) Oven Vanilla-greedy 2.44 2.66 3.13 2.86 3.35 3.35 3.48 3.64 Vanilla-sampling 2.45 2.65 3.10 2.84 3.33 3.37 3.45 3.61 VDD 2.20 2.44 2.96 2.67 3.10 3.21 3.39 3.59 OPERA 2.50 2.49 3.20 2.92 3.44 3.42 3.53 3.68 Woodpecker 2.46 2.46 3.21 2.90 3.45 3.45 3.52 3.70 LRV 2.51 2.52 3.18 2.88 3.47 3.40 3.57 3.74 LURE 2.49 2.51 3.18 2.91 3.42 3.46 3.55 3.72 PAI 2.54 2.60 3.26 2.95 3.48 3.48 3.60 3.78 HALC 2.52 2.57 3.24 2.94 3.50 3.45 3.61 3.80 VDGD (ours) 2.78 (+14%) 2.84 (+7%) 3.49 (+12%) 3.15 (+10%) 3.68 (+10%) 3.59 (+7%) 3.75 (+8%) 3.92 (+7%) LLaVA-Bench Vanilla-greedy 3.01 3.12 3.23 3.13 4.12 3.68 4.15 4.37 Vanilla-sampling 2.98 3.10 3.27 3.17 4.08 3.71 4.17 4.33 VDD 2.93 3.16 3.18 3.15 3.89 3.57 4.13 4.28 OPERA 3.08 3.18 3.30 3.22 4.16 3.75 4.22 4.40 Woodpecker 3.10 3.22 3.36 3.23 4.21 3.73 4.21 4.39 LRV 3.16 3.19 3.30 3.22 4.27 3.72 4.20 4.49 LURE 3.18 3.26 3.35 3.24 4.25 3.64 4.27 4.51 PAI 3.20 3.28 3.38 3.25 4.30 3.74 4.32 4.54 HALC 3.25 3.27 3.37 3.22 4.24 3.76 4.35 4.53 VDGD (ours) 3.36 (+12%) 3.47 (+11%) 3.52 (+9%) 3.38 (+8%) 4.45 (+8%) 3.96 (+8%) 4.50 (+8%) 4.78 (+9%) MMVET Vanilla-greedy 2.39 2.52 2.72 2.87 3.39 3.18 3.47 3.64 Vanilla-sampling 2.34 2.50 2.70 2.84 3.32 3.35 3.42 3.62 VDD 2.29 2.61 2.67 2.92 3.29 3.42 3.39 3.58 OPERA 2.48 2.67 2.77 3.01 3.37 3.39 3.52 3.69 Woodpecker 2.47 2.64 2.82 3.06 3.40 3.44 3.50 3.72 LRV 2.58 2.56 2.79 2.92 3.38 3.42 3.51 3.75 LURE 2.54 2.63 2.81 3.12 3.41 3.46 3.56 3.71 PAI 2.62 2.60 2.83 3.18 3.37 3.49 3.59 3.78 HALC 2.66 2.70 2.88 3.15 3.44 3.47 3.54 3.76 VDGD (ours) 2.85 (+19%) 3.01 (+19%) 3.23 (+19%) 3.45 (+20%) 3.78 (+12%) 3.57 (+12%) 3.74 (+8%) 3.99 (+10%) VaLLu Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82 3.23 3.45 Vanilla-sampling 1.86 2.01 1.64 2.18 2.70 2.83 3.21 3.40 VDD 1.47 1.55 1.80 1.80 2.32 2.63 3.19 3.42 OPERA 2.04 2.05 2.62 2.26 2.65 2.85 3.22 3.47 Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91 3.28 3.52 LRV 1.98 2.10 2.65 2.19 2.59 2.88 3.32 3.51 LURE 2.03 2.03 2.64 2.24 2.64 2.78 3.28 3.48 PAI 2.04 2.23 2.78 2.34 2.76 2.86 3.31 3.54 HALC 2.02 2.22 2.74 2.32 2.72 2.89 3.29 3.52 VDGD (ours) 2.16 (+11%) 2.64 (+30%) 3.16 (+20%) 2.72 (+24%) 3.45 (+30%) 3.01 (+7%) 3.48 (+8%) 3.67 (+6%)
Table 2: Performance comparison of VDGD with various baselines. VDGD outperforms by 2%-33%.
5.2 Experimental Setup
Datasets and Evaluation Metrics. For evaluation, we employ a variety of standard benchmarks focused on reasoning and information-seeking tasks. These include LLaVA-Bench, MM-Vet (Yu et al., 2023), MMBench (Liu et al., 2023d), MME (Fu et al., 2023), MathVista (test-mini subset), MathVision, and MMMU (validation set). For each benchmark, we employ our proposed GPT evaluation technique described in Section 3. Beyond being more robust to the diverse formats of model generation, we also find GPT scores to have a higher correlation to expert-human evaluation ($\approx$0.97; Section D has more details with original benchmark evaluation metrics that has a correlation of $\approx$0.92). We are inspired by a wealth of prior studies in hallucination evaluation (Sicong Leng et al., 2023; Yu et al., 2023; Liu et al., 2023c). All results are averaged across 3 runs.
Baselines. We compare VDGD against VCD, OPERA, Woodpecker, LRV, LURE, HALC (Chen et al., 2024a) and PAI (Liu et al., 2024d) hallucination mitigation techniques. Additionally, we compare two vanilla decoding methods (without any additional mitigation techniques): Vanilla-greedy with vanilla-greedy decoding and Vanilla-sampling with multinomial sampling-based decoding (top-p=0.5 and temperature=0.7). We employ greedy decoding for all methods as we find no difference in performance on sampling. We generate image descriptions for VDGD using the VCD mitigation technique. For LVLMs, we employ LLaVA-v1, LLaVA-v1.5, LLaVA-v1.6, mPLUG-Owl2, InternLM-X, CogVLM, Qwen2-VL (Wang et al., 2024b) and CogVLM2 (Hong et al., 2024).
Proposed VaLLu Benchmark. Automatic evaluation of LVLMs, with powerful closed-source models, on multiple large benchmarks can be prohibitively expensive. Thus, we propose VaLLu, a benchmark that we build by amalgamating existing benchmarks.
[…]
[trim= 0cm 1cm 0cm 0cm, width=]chart-11.pdf
@@ 5.2 Experimental Setup
VaLLu comprises a total of 1500 instances, with 554 sourced from Oven, 535 from MMMU, 218 from MMC, 74 from MathVista, 60 from HallusionBench, 53 from MATH-Vision, and 14 from MME. We only include instances with instructions that require open-ended generations and do not include any binary Yes/No or Multi-choice questions. We control the size to ensure the evaluation is affordable while keeping the diversity of tasks and topics for comprehensive analysis. VaLLu is manually filtered for noisy examples commonly found in existing benchmarks (we discuss this further in Appendix K.3), annotated with additional meta-data (illustrated in Fig. 9) and paired with an expert-provided response.
@@ 5.3 Results and Analysis
Quantitative results. Table 9 compares the performance of VDGD with other baselines. As we clearly see, VDGD not only outperforms all baselines but is also the only method that alleviates and improves performance on VaLLu. Results on other full benchmarks are provided in Appendix L.6, together with scores for other aspects beyond factuality, where we show that VDGD does not compromise the depth and expressivity of responses. Ablation results in Appendix C demonstrate that removing KL divergence leads to drop in performance. Additionally, using only the description prefix without the image leads to performance drops, confirming that VDGD enhances reasoning through grounding rather than making the model rely completely on the description alone.
Qualitative examples. Fig 10 compares responses from vanilla greedy decoding with VDGD. Grounding response generation results in more factual responses. Appendix B illustrates 2 failure cases of VDGD where incorrect descriptions can lead VDGD to hallucinate.
@@
[width=]vllm_sk.pdf
Figure 10: Illustration of responses generated with VDGD compared with greedy decoding. More in Sec. K.2.